Terri Ronsin had recently been transferred to the Home Security Systems Division of National Home Products. Shortly after taking over her new position as divisional controller, she was asked to develop the division’s predetermined overhead rate for the upcoming year. The accuracy of the rate is important because it is used throughout the year and any overapplied or underapplied over- head is closed out to Cost of Goods Sold at the end of the year. National Home Products uses direct labor-hours in all of its divisions as the allocation base for manufacturing overhead.

To compute the predetermined overhead rate, Terri divided her estimate of the total manufacturing overhead for the coming year by the production manager’s estimate of the total direct labor-hours for the coming year. She took her computations to the division’s general manager for approval but was quite surprised when he suggested a modification in the base. Her conversation with the general manager of the Home Security Systems Division, Harry Irving, went like this:

Ronsin: Here are my calculations for next year’s predetermined overhead rate. If you approve, we can enter the rate into the computer on January 1 and be up and running in the job-order costing system right away this year.

Irving:   Thanks for coming up with the calculations so quickly, and they look just fine. There is, how- ever, one slight modification I would like to see. Your estimate of the total direct labor-hours for the year is 440,000 hours. How about cutting that to about 420,000 hours?

Ronsin: I don’t know if I can do that. The production manager says she will need about 440,000 direct labor-hours to meet the sales projections for the year. Besides, there are going to be over 430,000 direct labor-hours during the current year and sales are projected to be higher next year.

Irving: Teri, I know all of that. I would still like to reduce the direct labor-hours in the base to some- thing like 420,000 hours. You probably don’t know that I had an agreement with your predecessor as divisional controller to shave 5% or so off the estimated direct labor-hours every year. That way, we kept a reserve that usually resulted in a big boost to net operating income at the end of the fiscal year in December. We called it our Christmas bonus. Corporate headquarters always seemed as pleased as punch that we could pull off such a miracle at the end of the year. This system has worked well for many years, and I don’t want to change it now.

Required:

Assume the following information:

Direct Materials

$40

per unit

Direct Labor

$20

per unit

Total Estimated Manufacturing Overhead

$8,400,000

Manufacturing overhead is allocated based on estimated direct-labor hours.

Each unit of product requires 1 direct labor hour.

  1. Calculate the cost of one unit of product, assuming that the overhead per unit is based on Terri Ronson’s estimate of 440,000 hours. (Round all dollar figures to two decimal places.)
    1. If 442,000 units were produced, how much overhead was applied to work in process.

  1. Calculate the cost of one unit of product, assuming that the overhead per unit is based on her supervisors preferred estimate of 420,000 hours. (Round all dollar figures to two decimal places.)
    1. If 442,000 units were produced, how much overhead was applied to work in process.

  1. During the year, the company produced and sold 442,000 units, and incurred actual overhead of $8,450,000, what is the under/overapplied overhead if:
    1. The estimated Direct Labor Hours is 440,000.
    2. The estimated Direct Labor Hours is 420,000.
    3. All over-applied and under-applied overhead applied directly to cost of goods sold. Assume that the company had $900,000 in net operating income before the over/under applied overhead adjustment is made. What is the revised net income after the over/underapplied overhead adjustment?
  2. Should Terri Ronson go along with the general manager’s request to reduce the direct labor hours in the predetermined overhead rate computation to 420,000 hours? Be sure to discuss the operational and ethical issues related to this decision.
440,000 Hours 420,000 Hours
Steps 1 and 2
Direct Materials $                40.00 $                40.00
Direct Labor $                20.00 $                20.00
Manufacturing Overhead
a.
b.
Product Cost Per Unit $                60.00 $                60.00
Overhead Applied (assuming 442,000 units) $                       -   $                       -  
Actual Overhead Incurred
Overhead Over/(Under) Applied $                       -  

$                       -   
Answer #1
Cost of unit product cost:
$
Direct materials 40
Direct labor 20
Overhead per unit (8400000/440000) 19.09
Unit product cost 79.09
Overhead applied to work in process=Units produced*overhead per unit=442000*19.09=$ 8437780
Cost of unit product cost:
$
Direct materials 40
Direct labor 20
Overhead per unit (8400000/420000) 20
Unit product cost 80
Overhead applied to work in process=Units produced*overhead per unit=442000*20=$ 8840000
Over/underapplied overhead if
The estimated Direct Labor Hours is 440,000
Actual overhead=8,450,000
Overhead applied= $ 8437780
Actual overhead > Overhead applied, Hence under application of overhead
Underapplied overhead=8450000-8437780=$ 12220
The estimated Direct Labor Hours is 420,000
Actual overhead=8,450,000
Overhead applied= $ 8840000
Actual overhead < Overhead applied, Hence over application of overhead
Overapplied overhead=8840000-8450000=$ 390000
Underapplied overhead debited to cost of goods sold.It will increase cost of goods sold and thus decrease net income
Revised net income=900000-12220=$ 887780
Overerapplied overhead credited to cost of goods sold.It will decrease cost of goods sold and thus increase net income
Revised net income=900000+390000=$ 1290000
Operationally it will be better to reduce direct labor hours since it will show that the workers are working at a higher efficiency.
It will enable the workers to get incentives
Ethically,It is not appropriate to reduce the direct labor hours since increased direct labor hours would be required due to increased demand
Manger should choose the ethical route.